OpenStudy (anonymous):

Sodium-24 has a half-life of 15 hours. How many grams of a 900 g sample will remain after 90 hours? 28.125 g 14.0625 g 7.5 g 7.03125 g

OpenStudy (unklerhaukus):

how many half-lives of Na-24 is 90 hours?

OpenStudy (anonymous):

15

OpenStudy (unklerhaukus):

one half life is 15 hours, how many half lives is 90 hrs?

OpenStudy (anonymous):

im really struggling with this sorry

OpenStudy (anonymous):

1350

OpenStudy (unklerhaukus):

every half life, the mass is reduced by half

OpenStudy (unklerhaukus):

the number of half lives is 90hr / 15 hr = ?

OpenStudy (anonymous):

so its 6

OpenStudy (unklerhaukus):

yes, 6 half lives, so divide the original mass in half, 6 times. ie 900/2^6

OpenStudy (anonymous):

14.0625

OpenStudy (unklerhaukus):

yep, [unit is grams]

OpenStudy (anonymous):

so is this my answer

OpenStudy (unklerhaukus):

does it make sense yet? initially the sample is 900g after 15hr : 450g ...30hr : 225g . . . 90hr : 14 g

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but can you help me with one more problem as well

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

A research study in 2000 found that there were 440 coyotes in a given region. The coyote population is expected to grow at a rate of 17% each year. In what year will the population first exceed 10,000?

OpenStudy (unklerhaukus):

year coyotes 2000 440 2001 440*114% 2002 440*114%*114% 200n 440*(114%)^n

OpenStudy (anonymous):

where did you get the 114

OpenStudy (unklerhaukus):

ops 100% + 17% = 117% (not 114%)

OpenStudy (anonymous):

ok so it would be 440(1+.17)^

OpenStudy (unklerhaukus):

yeah something like that/this \[N(t) = 440\times(117\%)^t\]

OpenStudy (unklerhaukus):

What value of t makes N = 1000

OpenStudy (unklerhaukus):

ie, set N(t) = 1000, and solve for t

OpenStudy (anonymous):

wait can you explain the 1000 part

OpenStudy (unklerhaukus):

[damnit, i cant read very well today] The question say `In what year will the population first exceed 10,000?` so set N(t) = 10,000

OpenStudy (anonymous):

so do i multiply to 10,000 by a number

OpenStudy (unklerhaukus):

\[10,000=440(117\%)^t\]divide both sides by 440, then . . .

OpenStudy (anonymous):

this is the part i didnt understand

OpenStudy (anonymous):

do i divide 10,000 by 440 and by 1.17

OpenStudy (unklerhaukus):

\[10,000=440(117\%)^t\\ \frac{10,000}{440}=(117\%)^t\]

OpenStudy (unklerhaukus):

now you have an exponential, to simplify this , use logs

OpenStudy (anonymous):

i got 22

OpenStudy (anonymous):

so does t=22

OpenStudy (anonymous):

so it would take 22 years

OpenStudy (unklerhaukus):

\[10,000=440(117\%)^t\\ \frac{10,000}{440}=(117\%)^t\\ 22.7 = (117\%)^t\\ 22.7 = 1.17^{t}\\ \log_{1.17}22.7 = \log_{1.17}1.17^{t}\\[2ex] \log_{1.17}22.7=t\]

OpenStudy (unklerhaukus):

now use the change of base formula \[\log_ba=\frac{\ln a}{\ln b}\]

OpenStudy (anonymous):

so the year would be 2022

OpenStudy (unklerhaukus):

how did you get 22?

OpenStudy (anonymous):

i dont know

OpenStudy (unklerhaukus):

[well its not correct]

OpenStudy (unklerhaukus):

what do you get when you simplify \[\log_{1.17}22.7\]?

OpenStudy (unklerhaukus):

(use a calculator)